OpenStudy (anonymous):

how do I find the zeros (roots) of x^2+9

OpenStudy (campbell_st):

well they will be complex you are solving an equation... so you need to let \[x^2 + 9 = 0\] you need to find the values of x that make the equation zero... and there will be 2 zeros as its a quadratic equation. subtract 9 from both sides of the equation and then take the square root of both sides. one solution will be positive and the other negative hope it helps

OpenStudy (anonymous):

so like x+3 and x-3?

OpenStudy (campbell_st):

not quite... do it step by step... subtract 9 from both sides of the equation... what do you get?

OpenStudy (anonymous):

x=-9

OpenStudy (anonymous):

well x^2=-9

OpenStudy (campbell_st):

nearly \[x^2 = -9\] now take the square root of both sides..

OpenStudy (anonymous):

it would be x and the sqr root 3 i?

OpenStudy (anonymous):

so is it root 3 i?

OpenStudy (anonymous):

positive and negative

OpenStudy (campbell_st):

well nearly \[x= \pm \sqrt{-9} ~~~or~~~ x = \pm \sqrt{9i^2}\] so hopefully this will help you get the answer...

OpenStudy (anonymous):

sweet thanks

OpenStudy (campbell_st):

you don't need the root since 9 is a square number

OpenStudy (anonymous):

alrighty thanks dude